A 0.109-kg particle undergoes simple harmonic motion along the horizontal x-axis between the points ㄨㄧ--0.279 m and X2 = 0.499 m. The period of oscillation is 0.557 s. Find the frequency, the equilibrium position, Xeq, the amplitude, A, the maximum speed, Vmax, the maximum magnitude of acceleration, amax, the force constant, k, and the total mechanical energy, Etot
Answer #1

Feel free to ask any doubt regarding this question...

Solution 厅- -0.279 ostion 1395 Hette xz o. meter A- meter 0. 389 meter(VI) k = 13.868 시나 2

Stilif you have any doubt than ask me in the comments section

Rate this answer....thanks
Answer #2

The frequency is the inverse of the period, T.

f=1T

Substitute the given value.

f=10.685 s=1.46 Hz

The equilibrium position is the midpoint of the two endpoints. Its coordinate is

xeq=x1+x22

Substitute the data.

xeq=(0.253 m)+(0.471 m)2=0.109 m

The amplitude is half the distance between the endpoints.

A=|x1x2|2

Its value is

A=|(0.253 m)(0.471 m)|2=0.362 m

The maximum speed is given by

vmax=2πfA=π|x1x2|T

vmax=π×|(0.253 m)(0.471 m)|0.685 s=3.32 m/s

The maximum magnitude of acceleration is

amax=(2πf)2A=2|x1x2|(πT)2

Its value is

amax=2×|(0.253 m)(0.471 m)|(π0.685 s)2=30.5 m/s2

To find the force constant, start with the formula for the period,

T=2πmk

where m denotes the mass. Now solve for the force constant.

k=m(2πT)2

Insert the given values.

k=(0.105 kg)(2π0.685 s)2=8.83 N/m

To find the total mechanical energy, use the fact that it is conserved and evaluate it at any point along the motion. At the equilibrium position, the mechanical energy is completely kinetic, with the particle moving at its maximum speed, vmax, which you found previously. The total mechanical energy, therefore, equals the maximum kinetic energy.

Etot=12mvmax2=m2(π|x1x2|T)2

Substitute the given values.




answered by: Muhammad Aslam